# Intentional Gesture: Deliver Your Intentions with Gestures for Speech

### v1

@@ Abstract
When humans speak, gestures help convey communicative intentions, such as adding emphasis or describing concepts. However, current co-speech gesture generation methods rely solely on superficial linguistic cues (e.g. speech audio or text transcripts), neglecting to understand and leverage the communicative intention that underpins human gestures. This results in outputs that are rhythmically synchronized with speech but are semantically shallow. To address this gap, we introduce Intentional-Gesture, a novel framework that casts gesture generation as an intention-reasoning task grounded in high-level communicative functions. First, we curate the InG dataset by augmenting BEAT-2 with gesture-intention annotations (i.e., text sentences summarizing intentions), which are automatically annotated using large vision-language models. Next, we introduce the Intentional Gesture Motion Tokenizer to leverage these intention annotations. It injects high-level communicative functions (e.g., intentions) into tokenized motion representations to enable intention-aware gesture synthesis that are both temporally aligned and semantically meaningful, achieving new state-of-the-art performance on the BEAT-2 benchmark. Our framework offers a modular foundation for expressive gesture generation in digital humans and embodied AI. Project Page: https://andypinxinliu.github.io/Intentional-Gesture
@@ 1 Introduction
Gestures accompany speech to convey emotions, emphasize ideas, and facilitate comprehension [6, 30, 3], forming an essential part of human communication. As AI advances, enabling virtual avatars to produce expressive gestures will be critical for digital avatar construction [32, 29, 9, 27, 28].
Recent works [35, 16, 13, 38] have explored full-body gesture generation conditioned on speech audio or text transcripts. However, these approaches largely treat raw linguistic input as sufficient semantics, overlooking the deeper communicative intentions—such as emphasis, deixis, and affirmation—that implicitly govern gesture behavior. Recovering these latent functions is essential for generating semantically coherent non-verbal expressions.
Early systems [19, 26, 20] attempted to derive these deeper communicative functions through rule-based templates, but such approaches lack scalability and domain adaptability. Inspired by the linguistic–gesture link in psycholinguistics, we reframe gesture modeling as a reasoning task: the model first understands the reasons or communicative intentions behind the speech, and then treats gestures as their downstream realizations.
To achieve this, we first augment the BEAT-2 dataset [14] with structured annotations of inferred communicative functions via VLMs (Vision-Language-Models), creating a large-scale InG (Intention-Grounded) dataset. This enriched corpus links speech, inferred intentions, and corresponding gesture realizations, enabling intention-aware gesture generation for the first time.
@@
Building upon this understanding model, we introduce the Intentional Gesture Tokenizer, a novel quantization model that directly embeds intention semantics into the motion representation space. Unlike prior works that discretize body parts independently without semantic grounding [21, 15], our tokenizer processes global body motion holistically and supervises latent representations using intention-aware motion features. This design ensures that discrete motion tokens encode not only fine-grained motion patterns but also communicative meaning. With the proposed tokenizer, our approach generates gestures that are not only temporally synchronized but also semantically expressive and interpretable, advancing human-avatar communication. Our contributions can be summarized as follows:
@@ 2 Related Works
Co-speech Gesture Generation. Existing works on co-speech gesture generation typically use skeleton- or joint-level pose representations. Several methods [17, 7, 36, 12, 42, 16] learn hierarchical semantics or apply contrastive learning to align audio and gesture embeddings. HA2G [17] constructs multi-level audio-motion embeddings, while TalkShow [38], CaMN [14], and EMAGE [13] introduce large-scale datasets for joint face-body modeling with GPT-style decoding. More recent models such as MambaTalk [35], DiffSHEG [5], and GestureLSM [15] focus on efficient flow matching and spatiotemporal modeling for quality-speed tradeoffs. Semantic Gesticulator [43] and RAG-Gesture [21] leverage LLMs to retrieve discourse-relevant gestures as references. In contrast, we explore the modeling of communicative intentions as explicit semantics to control gesture generation.
Vision Tokenization for Generation. In visual generation, tokenization encodes raw pixels into compact representations [34, 25]. Vector-quantization [34] enable discrete latent spaces compatible with autoregressive generation [33, 31, 4, 40]. Recent studies like REPA [41] and Re.vs.Gen [37] show that aligning generation representation with understanding improves synthesis quality. We further introduce the Intentional Gesture Tokenizer, supervising motion tokenization with intention-aligned semantics, and demonstrate this strategy significantly enhances gesture generation quality.
Figure 2: Overview of the annotation pipeline. Sentence-level segments and word-aligned keyframes are provided to the VLM, which generates (1) motion descriptions across body regions, (2) communicative function labels, (3) gesture behavior mappings, and (4) inferred intentions. This structured annotation enables scalable creation of a semantically grounded, visually aligned gesture dataset.
@@ 3 Intentional Gesture Dataset (InG)
Gestures systematically reflect semantic, pragmatic, and rhetorical structures [19, 20, 26]; for example, affirmation to nodding and comparison to lateral gestures aligned with spatial metaphors. Modeling these communicative functions is critical for generating expressive gestures. Marsella et al. [19] introduced a rule-based system that derives communicative functions from syntactic and lexical patterns, mapping them to gesture classes via hand-crafted rules. While effective, such systems rely on fixed rules and limited dictionaries, restricting scalability and domain coverage. Building on these insights, we propose a scalable pipeline using VLMs to automatically infer communicative functions and map them to gestures, enabling large-scale, data-driven semantic gesture modeling.
@@ 3.1 Dataset Construction
We introduce InG (Intention-Grounded Gestures), a dataset that augments BEAT-2 [14] with structured, intention-aware annotations linking linguistic functions to gesture behaviors. Unlike prior datasets focused on motion fidelity or prosody alignment, InG enables intention-controllable gesture generation. We employ a two-stage VLM-based strategy, illustrated in Fig. 2.
Annotation Protocol. For training data, the VLM receives sentence context and word-aligned keyframes, allowing inference of communicative functions, motion descriptions, and speaker intent. For testing data, annotations are generated from transcripts only to prevent leakage from observed gestures. Detailed prompting strategies and validations are provided in the Appendix.
Sentence-Level Segmentation and Keyframe Extraction. To prepare BEAT-2 for intention annotation, we render each sequence into frame-wise visual data. Speech sequences are segmented into sentence-level units using Whisper-X [1], which restores punctuation and provides word-level alignment. Boundaries are identified via punctuation, and timestamps are extracted. For each word, we uniformly sample one keyframe within its temporal span to capture the pose at word-level granularity.
Prompt-Based Gesture Intention Annotation. Given the segmented utterances and word-level keyframes, we design a structured prompt to guide GPT-4o-mini in multi-stage annotation: (1) Motion Analysis: describing body poses across regions (head, arms, fingers, torso); (2) Function Derivation: identifying communicative functions (e.g., Emphasis, Negation, Deixis); (3) Gesture Behavior Mapping: associating functions with gesture forms; (4) Inferred Intention: synthesizing how verbal and nonverbal cues reflect speaker goals. This structured approach enables scalable creation of semantically and physically grounded gesture annotations.
@@ 3.2 Dataset Statistics and Analysis
While function derivations follow a well-defined ontology, gesture behavior mappings and inferred intentions exhibit open-ended variability across speakers and contexts. To ensure interpretability, our prompt design grounds mappings in established literature (McNeill [20], Kendon [26]). Additional statistics, qualitative examples and function co-occurrence analysis are provided in Appendix.
@@
Final preferences are determined by majority vote. Shown in Fig. 4, both VLM protocols outperform human annotations, especially on Intent Alignment and Overall Preference. Inter-rater agreement averaged 0.76 Fleiss’ $\kappa$, indicating substantial consistency. These results validate the effectiveness of our structured prompting and the scalability of VLM-based annotation. Additional validation details, including hallucination audits and mapping coherence analysis, are in Appendix A.3–A.6.
@@ 4.1 Intention Understanding
To leverage textual intention annotations for co-speech gesture synthesis (akin to text-to-image generation), we propose a CLIP-like understanding model, H-AuMoCLIP, for gestures to better encode textual intention features. We build this Intention-Audio-Motion CLIP on top of TANGO’s [12] audio-motion alignment model (AuMoCLIP) by explicitly incorporating communicative intent.
Intentional Semantic Fusion. AuMoCLIP trains its joint embedding space using a two-tower contrastive learning framework, where its audio tower typically combines acoustic features (e.g., wav2vec2) with BERT-encoded speech transcripts. Building upon this, we introduce a dedicated pathway, Intentional Semantic Fusion, for merging intention semantics to create a richer, intention-aware joint embedding space. Our Intentional Semantic Fusion mechanism fuse transcript and intention embeddings via a linguistically grounded mechanism. Following TANGO, we align transcript tokens to the audio timeline using a CTC-based model, and encode both the transcript and annotated intent with separate BERT encoders. In addition to TANGO, the aligned transcript embeddings serve as queries in a cross-attention module, with intention embeddings as keys and values. The output captures context-specific communicative cues, which are then concatenated with wav2vec2 audio features to form the final high-level representation. This composite embedding is used in contrastive training to jointly align audio, motion, and intent. Note that we retain TANGO’s low-level contrastive training setup to align audio-motion features rhythmic level as well.
@@ 4.2 Intentional Gesture Tokenizer
Prior works [21, 15, 13, 35] often model different body regions using separate quantizers, maintaining distinct encoders, decoders, and codebooks per body part. While this tokenizer encourages local disentanglement, it introduces three limitations: (1) high training and inference cost, (2) poor global body coherence, and (3) weak connection to speech-level semantics. To address these issues, we propose an Intentional Gesture Tokenizer, which introduces two core innovations: (1) learning disentangled latent factors within a unified global body representation through multi-codebook quantization, and (2) embedding explicit intention semantics into the motion representation through semantic supervision.
Architecture Design. While CNNs are effective at capturing fine-grained local details [23, 2, 15, 13], Transformers offer greater scalability with reduced inductive bias [8]. To leverage the strengths of both, we adopt a hybrid VQ-tokenizer architecture: the encoder comprises CNN blocks that downsample the input gesture motion by a factor of 4, followed by 1 Transformer layers and a vector quantizer that outputs discrete latent codes. The decoder mirrors this structure in reverse: Transformer layers followed by CNN blocks that upsample the features to reconstruct the gesture motion.
@@ 4.3 Gesture Generation
We then leverage Sec.4.1 and Sec.4.2 to enhance gesture generation. We adapt a state-of-the-art gesture generation model, GestureLSM [15], by replacing its original audio encoder and motion tokenizer. Specifically, we replace GestureLSM’s audio encoder with our intent-aware audio encoder, which extracts hierarchical representations capturing both rhythmic audio features and communicative intentions. Furthermore, we replace its RVQ-based tokenizer with our Intentional Gesture Tokenizer. Additional architectural details of the original GestureLSM are deferred to the Appendix.
@@ 5 Experiments
We conduct all experiments on BEAT2 [13], comprises 60 hours of high-quality SMPL-based gesture data collected from 25 speakers (12 female, 13 male). The dataset contains 1,762 sequences, each with an average duration of 65.66 seconds, where each sequence captures responses to daily inquiries. For consistency, we adopt the train-validation-test split protocol defined in EMAGE [13].
@@ 5.1 Quantitative Comparisons
Metrics. We evaluate Fréchet Gesture Distance (FGD)[39] for distributional similarity between ground truth and synthesized gestures. Diversity (Div.)[10] is measured by the average L1 distance across gesture clips, while speech-motion synchronization is assessed via Beat Constancy (BC) [11].
Evaluation Results. We evaluate both quality and generalization in three settings: single-speaker, all-speaker, and zero-shot (training on 20 speakers and testing on 5 unseen speakers). As shown in Table 1, our method significantly outperforms baselines across all settings. These gains are attributed to our intentional motion representation and intention control, which reduce unnatural gesture patterns. More detailed comparisons are provided in the Appendix. To validate runtime efficiency, we generate intention descriptions from transcripts via an API prior to model execution. Our method retains the efficiency of GestureLSM [15], while improving performance through a unified motion representation, and achieves real-time generation at 30.4 FPS on an H100 GPU.
Figure 6: The subjective Comparisons. Compared with other methods, Intentional-Gesture presents more natural and coherent motion patterns to represent specific words or phrases (highlighted in red).

### v2

# Intentional-Gesture: Deliver your Intentions with Gestures for Speech

@@ Abstract
When humans speak, gestures help convey communicative intentions, such as adding emphasis or describing concepts. However, current co-speech gesture generation methods rely solely on superficial linguistic cues (e.g. speech audio or text transcripts), neglecting to understand and leverage the communicative intention that underpins human gestures. This results in outputs that are rhythmically synchronized with speech but are semantically shallow. To address this gap, we introduce Intentional-Gesture, a novel framework that casts gesture generation as an intention-reasoning task grounded in high-level communicative functions. First, we curate the InG dataset by augmenting BEAT-2 with gesture-intention annotations (i.e., text sentences summarizing intentions), which are automatically annotated using large vision-language models. Next, we introduce the Intentional Gesture Motion Tokenizer to leverage these intention annotations. It injects high-level communicative functions (e.g., intentions) into tokenized motion representations to enable intention-aware gesture synthesis that are both temporally aligned and semantically meaningful, achieving new state-of-the-art performance on the BEAT-2 benchmark. Our framework offers a modular foundation for expressive gesture generation in digital humans and embodied AI.
@@ 1 Introduction
Gestures accompany speech to convey ideas and facilitate comprehension De Ruiter et al. (2012); Song et al. (2023); Burgoon et al. (1990), forming an essential part of human communication. As AI advances Lu et al. (2023; 2024b); Ren et al. (2025), enabling virtual avatars to produce expressive gestures will be critical for digital avatar construction Tang et al. (2025); Song et al. (2024c); Huang et al. (2024); Song et al. (2024a; b).
Recent works Liu et al. (2025b; 2023); Yi et al. (2023) have explored full-body gesture generation conditioned on speech audio or text transcripts. However, these approaches largely treat raw linguistic input as sufficient semantics, overlooking the deeper communicative intentions—such as emphasis, deixis, and affirmation—that implicitly govern gesture behavior. Recovering these latent functions is essential for generating semantically coherent non-verbal expressions.
Early systems Marsella et al. (2013); Saund & Marsella (2021); McNeill (1992) attempted to derive these deeper communicative functions through rule-based templates, but such approaches lack scalability. Inspired by the linguistic–gesture link in psycholinguistics, we reframe gesture modeling as a reasoning task: the model first understands the reasons or communicative intentions behind the speech, and then treats gestures as their downstream realizations.
To achieve this, we first augment the BEAT-2 dataset Liu et al. (2022b) with structured annotations of inferred communicative functions via VLMs (Vision-Language-Models), creating a large-scale InG (Intention-Grounded) dataset. This enriched corpus links speech, inferred intentions, and corresponding gesture realizations, enabling intention-aware gesture generation for the first time.
@@
Building upon this understanding model, we introduce the Intentional Gesture Tokenizer, a novel quantization model that directly embeds intention semantics into the motion representation space. Unlike prior works that discretize body parts independently without semantic grounding Mughal et al. (2025); Liu et al. (2025a), our tokenizer processes global body motion holistically and supervises latent representations using intention-aware motion features. This design ensures that discrete motion tokens encode not only fine-grained motion patterns but also communicative meaning. With the proposed tokenizer, our approach generates gestures that are not only temporally synchronized but also semantically expressive and interpretable, advancing human-avatar communication. Our contributions can be summarized as follows:
@@ 2 Related Works
Co-speech Gesture Generation. Existing works on co-speech gesture generation typically use skeleton- or joint-level pose representations. Several methods Liu et al. (2022c); Deichler et al. (2023); Xu et al. (2023); Liu et al. (2024a); Zhang et al. (2024a); Liu et al. (2025b) learn hierarchical semantics or apply contrastive learning to align audio and gesture embeddings. HA2G Liu et al. (2022c) constructs multi-level audio-motion embeddings, while TalkShow Yi et al. (2023), CaMN Liu et al. (2022b), and EMAGE Liu et al. (2023) introduce large-scale datasets for joint face-body modeling with GPT-style decoding. More recent models such as MambaTalk Xu et al. (2024), DiffSHEG Chen et al. (2024b), and GestureLSM Liu et al. (2025a) focus on efficient flow matching Li et al. (2025); Zhu et al. (2024) and spatiotemporal modeling. Semantic Gesticulator Zhang et al. (2024b) and RAG-Gesture Mughal et al. (2025) leverage LLMs to retrieve discourse-relevant gestures as references. In contrast, we explore the modeling of communicative intentions as explicit semantics to control gesture generation.
Vision Tokenization for Generation. In visual generation Lu et al. (2025); Gao et al. (2024); Lu et al. (2024a), tokenization encodes raw pixels into compact representations Van Den Oord et al. (2017); Rombach et al. (2022). Vector-quantization Van Den Oord et al. (2017) enable discrete latent spaces compatible with autoregressive generation Tian et al. (2024); Sun et al. (2024); Chang et al. (2022); Yu et al. (2023). Recent studies like REPA Yu et al. (2024) and Re.vs.Gen Yao et al. (2025) show that aligning generation representation with understanding improves synthesis quality. We further introduce the Intentional Gesture Tokenizer, supervising motion tokenization with intention-aligned semantics, and demonstrate this strategy significantly enhances gesture generation quality.
Figure 2: Overview of the annotation pipeline. Sentence-level segments and word-aligned keyframes together with rule-based descriptions are provided to the VLM, which generates (1) motion descriptions across body regions, (2) communicative function labels, (3) gesture behavior mappings, and (4) inferred intentions. This structured annotation enables scalable creation of a semantically grounded, visually aligned gesture dataset.
@@ 3 Intentional Gesture Dataset (InG)
Gestures systematically reflect semantic, pragmatic, and rhetorical structures Marsella et al. (2013); McNeill (1992); Saund & Marsella (2021); for example, affirmation to nodding and comparison to lateral gestures aligned with spatial metaphors. Modeling these communicative functions is critical for generating expressive gestures.  Marsella et al. (2013) introduced a rule-based system that derives communicative functions from syntactic and lexical patterns, mapping them to gesture classes via hand-crafted rules. While effective, such systems rely on fixed rules and limited dictionaries, restricting scalability and domain coverage. Building on these insights, we propose a scalable pipeline using VLMs to automatically infer communicative functions and map them to gestures, enabling large-scale, data-driven semantic gesture modeling.
@@ 3.1 Dataset Construction
We introduce InG (Intention-Grounded Gestures), a dataset that augments BEAT-2 Liu et al. (2022b) and Audio2PhotoReal Ng et al. (2024) with structured, intention-aware annotations linking linguistic functions to gesture behaviors. Unlike prior datasets focused on motion fidelity or prosody alignment, InG enables intention-controllable gesture generation. We employ a two-stage VLM-based strategy, illustrated in Fig. 2.
Annotation Protocol. We design a unified annotation pipeline that integrates interpretable rule-based motion descriptors with structured VLM prompting to yield intention-aware gesture annotations. For training data, the VLM receives sentence context together with word-aligned keyframes, enabling inference of communicative functions, motion descriptions, and speaker intent. For testing data, annotations are generated solely from transcripts to prevent leakage from observed gestures. Detailed prompting strategies and validation procedures are provided in the Appendix.
Rule-Based Motion Description Generation. To provide interpretable motion descriptors aligned with linguistic units, we design a rule-based analysis procedure that converts raw pose sequences into structured gesture descriptions. The process operates in three stages: (1) Temporal Windowing. Each sentence is divided into sub-windows of 1–2 seconds based on word-level timestamps. Consecutive words within the same span are grouped together, while longer segments are recursively partitioned. This ensures that each window corresponds to a manageable co-speech gesture unit. (2) Canonicalized Motion Representation. All sequences are normalized to the canonical space relative to the body root. For elbows, wrists, head, and fingers, motions are represented by joint pose angles, while for hands, we track 3D positional displacements over time. (3) Movement Characterization. Within each window, trajectories are smoothed and segmented by direction and amplitude. Segments are categorized into qualitative tiers (very slight, slight, moderate, significant), where thresholds are defined relative to the standard deviation of motion distributions and refined through pilot annotation experiments and human evaluation. Detected patterns are then labeled as monotonic (e.g., “slightly forward”), bidirectional (e.g., “forward then backward”), or oscillatory (e.g., “repeated side-to-side”). Each description is anchored to visual evidence by presenting the first, last, and a representative keyframe from the corresponding word span, ensuring that textual descriptors remain verifiable against pose snapshots. We defer algorithm details in the Appendix.
Prompt-Based Gesture Intention Annotation. The rule-based descriptors and reference frames form the grounding for higher-level annotation with GPT-4o-mini. Given segmented utterances and keyframes, we design a structured prompt for multi-stage annotation: (1) Motion Analysis — describing body poses across regions (head, arms, fingers, torso) using rule-based descriptors together with reference frames; (2) Function Derivation — identifying communicative functions (e.g., emphasis, negation, deixis); (3) Gesture Behavior Mapping — linking communicative functions to gesture forms; and (4) Inferred Intention — synthesizing how verbal and nonverbal cues reflect speaker goals. This layered approach ensures that gesture annotations are both semantically interpretable and physically grounded, enabling intention-controllable gesture generation.
Human-in-the-Loop Validation. To ensure reliability of the VLM-generated annotations, we adopt a human-in-the-loop filtering stage. For each utterance, the LLM is instructed to produce five independent candidate annotations following the structured protocol described above. These candidates vary in phrasing and granularity, but all conform to the same motion–function–intention schema. Human labelers then evaluate the set of five responses and remove those that contain incorrect interpretations, inconsistent mappings, or implausible intentions. Only the remaining high-quality annotations are retained in the dataset.
@@ 3.2 Dataset Statistics and Analysis
While function derivations follow a well-defined ontology, gesture behavior mappings and inferred intentions exhibit open-ended variability across speakers and contexts. To ensure interpretability, our prompt design grounds mappings in established literature (McNeill McNeill (1992), Kendon Saund & Marsella (2021)). Additional statistics and analysis are provided in Appendix.
@@
Final preferences are determined by majority vote. Shown in Fig. 4, both VLM protocols outperform human annotations. Inter-rater agreement averaged 0.76 Fleiss’ $\kappa$, indicating substantial consistency.
@@ 4.1 Intention Understanding
To leverage textual intention for co-speech gesture synthesis, we propose a CLIP-like understanding model, H-AuMoCLIP, for gestures to better encode textual intention features. We build this Intention-Audio-Motion CLIP on top of TANGO’s (Liu et al., 2024a) audio-motion alignment model (AuMoCLIP) by explicitly incorporating communicative intent.
Intentional Semantic Fusion. AuMoCLIP trains its joint embedding space using contrastive learning, where its audio tower combines acoustic features with BERT-encoded speech transcripts. Building upon this, we introduce Intentional Semantic Fusion for merging intention semantics to create a richer, intention-aware joint embedding space. Our Intentional Semantic Fusion mechanism fuse transcript and intention embeddings via a linguistically grounded mechanism. Following TANGO, we align transcript tokens to the audio timeline using a CTC-based model, and encode both the transcript and annotated intent with separate BERT encoders. In addition to TANGO, the aligned transcript embeddings serve as queries in a cross-attention module, with intention embeddings as keys and values. The output captures context-specific communicative cues, which are then concatenated with wav2vec2 audio features to form the final high-level representation. This composite embedding is used in contrastive training to jointly align audio, motion, and intent.
@@ 4.2 Intentional Gesture Tokenizer
Prior works Mughal et al. (2025); Liu et al. (2025a; 2023) model different body regions using separate encoders, decoders, and codebooks per body part. While this design encourages local disentanglement, it has three limitations: (1) high training and inference cost, (2) poor global body coherence, and (3) weak connection to speech-level semantics. To address these issues, we propose an Intentional Gesture Tokenizer with two core innovations: (1) learning disentangled latent factors within a unified global body representation through multi-codebook quantization, and (2) embedding explicit intention semantics into the motion representation through semantic supervision.
@@ 4.3 Gesture Generation
We then leverage Sec.4.1 and Sec.4.2 to enhance gesture generation. We adapt GestureLSM (Liu et al., 2025a), by replacing its original audio encoder and motion tokenizer. Specifically, we replace GestureLSM’s audio encoder with our intent-aware audio encoder, which extracts hierarchical representations capturing both rhythmic audio features and communicative intentions. Furthermore, we replace its RVQ-based tokenizer with our Intentional Gesture Tokenizer. Additional architectural details of the original GestureLSM are deferred to the Appendix.
@@ 5 Experiments
We conduct main experiments on BEAT2 Liu et al. (2023), which comprises 60 hours of high-quality SMPL-based conversational or speach gesture data collected from 25 speakers. The dataset contains 1,762 sequences, each with an average duration of 65.66 seconds, following the train-validation-test split protocol defined in EMAGE Liu et al. (2023). In addition, we also explores its application on Audio2Photoreal Ng et al. (2024), which provides about 8 hours of dyadic interactions between listening and speaking actions. We defer the experiment results on Audio2PhotoReal to Appendix with further video demos provided in the supplementary material.
@@ 5.1 Quantitative Comparisons
Metrics. We evaluate Fréchet Gesture Distance (FGD) Yoon et al. (2020) for pose sequence angle distributional similarity, Diversity (Div.) Li et al. (2021a) as the average L1 distance across clips, and Beat Constancy (BC) Li et al. (2021b) for speech-motion synchronization.
Evaluation Results. We evaluate both quality and generalization in three settings: single-speaker, all-speaker, and zero-shot (training on 20 speakers and testing on 5 unseen speakers). As shown in Tab. 1, our method significantly outperforms baselines across all settings. These gains are attributed to our intentional motion representation and intention control, which reduce unnatural gesture patterns. More detailed comparisons are provided in the Appendix. To validate runtime efficiency, we generate intention descriptions from transcripts via an API prior to model execution. Our method retains the efficiency of GestureLSM Liu et al. (2025a), while improving performance through a unified motion representation, and achieves generation at 30.4 FPS on an H100 GPU.
Figure 6: The subjective Comparisons. Compared with others, Intentional-Gesture presents more natural and coherent motion patterns to represent specific words or phrases (highlighted in red).